Shift report 20011110 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Nov 10 2001 - 00:01:36 EST

  • Next message: Apache: "Shift report 20011110 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Nov 10 00:01:36 2001, Shift: 16:00-24:00
    Supervisor: BHS, with: 
    Log book#10 pages 119 - 121
    
    Shift Summary:
    ------------------------------------------------------------
    1: Had a brief store at the start of the shift, then STAR 
    broke something and needed an IR access. This is now over
    and MCR is promising an injection soon.
    
    2: One data run:
    5743   210k trig4, 65k trig6
           MRS@90, D5@0 field
           FFS@20, D1-D4@1/10B (BFS@6, D3 off)
    
    Polarity of D5 was then switched to A.
    
    3: D5 plane 1 tripped once.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Nov 10 2001 - 00:02:19 EST